JeliseWh05:

Can someone help me with this?? “summing of three cubes”: Find x, y, and z such that x³+y³+z³=k, for each k from 1 to 100.

jhonyy9:

1^3 = 1 2^3 = 8 3^3 = 27 4^3 = 64 so from these cubed numbers assuming 3 how can you make 100 ? not possible bc. then you assume the 3 greatest you get 99 and then you write the next cubed number so 5^3 = 125 so now then you wan to play with these cubed numbers you can check how much sum can you getting for k ? hope helped

imqwerty:

No constraints on x,y, and z?